Carla is 5 years old. Pete's age is the sum of Carla and Tyler's age. Tyler is 13 years younger than Pete. How old are Pete and Tyler?
Answer by TimothyLamb(4379)

x = carla age now
y = pete age now
z = tyler age now
---
x = 5
y = x + z
z = y - 13
---
put the system of linear equations into standard form
---
x = 5
x - y + z = 0
y - z = 13
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
This linear system is singular. It has no solution
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php